OpenStudy (anonymous):

If logM/N=5 and logP/N=7, what can you say about the relationship between M and P? A. M=2P B. P=2M C. P=100M D. P=10M

OpenStudy (anonymous):

I assume you mean log(M/N)? log (P/N)?

OpenStudy (anonymous):

yes \[\log \frac{ M }{ N } and \log \frac{ P }{ N }\]

OpenStudy (anonymous):

log (M/N) = log M - log N, so log M - log N = 5 log (P/N) = log P - log N, so log P - log N = 7 If you subtract equation 2 from equation 1, you have log M - log P = -2 That means that log (M/P) = -2 ----> 10^-2 = M/P ---> P = 100M

OpenStudy (anonymous):

thank you SO much! :)

OpenStudy (anonymous):

@ali1029 Does each step make sense?

OpenStudy (anonymous):

yeah I see exactly what you did! it seems kind of like you subtracted the rate of change or whatever it is called (I dont know exact terms :p )

OpenStudy (anonymous):

but yeah i see what you did, thanks!! :)

OpenStudy (anonymous):

Welcome. (Rate of change is the slope...but slope was not considered in this problem..you probably meant something else).